Income    Statement (A-H)              

                       

Data for current year              

                       

Number of X units sold          14,000

Sales price per unit                 1,200

Cost per unit                           300

Commission (% of sales)        10%

Freight cost per unit                200

                       

Number of Y units sold          19,000

Sales price per unit                 14,800

Cost per unit                           600

Commission (% of sales)        20%

Freight cost per unit                200

Advertising per unit                400

                       

Travel expense                        180,000

Marketing expense                  77,000

Utility expense                        15,000

Salaries                                   543,000

Salary overhead                      30%

                       

Cost of Machine M                 1,485,000

Years to depreciate                 9

Cost of Machine N                  656,000

Years to depreciate                 4

                       

Tax rate                                               40%

                       

Dividend Payout Ratio                       20%

                                               

Data just prior to current year            

                       

Paid In Capital                                                13,000,000

Additional Paid in Capital                  128,000,000

Retained Earnings                               149,000,000

                       

                       

  1. What was the Sales?                       

                       

                       

      2. What was the Net Income?            

                       

   3. What is the Return on Equity?                       
Answer #1

1. Sales of company( A-H)

Sales = units sold * unit price

product X = 14000*1200 = 1,68,00,000

product Y = 19800* 14800 = 28,12,00,000

Total sales = 29,80,00,000

2. Computation of Net income of ( A-H)

Total revenue    = 29,80,00,000

Less :Cost of goods sold

PRODUCT   X -  14000*300 = (42,00,000)

PRODUCT Y -  19000*600 = (1,14,00,000)

Gross profit = 28,24,00,000

Less Expenses:

Salaries = ( 543,000)

Salary overhead   = ( 162900) i.e = 543000*30%

Travel expense = (180,000)

Marketing expense = (77,000)

Utility expense = (15,000)

Freight cost   

product X 200*14000 = (28,00,000)

product Y 200*19000   = (38,00,000)

Advertising Cost (product Y) 400 * 19000 = (76,00,000)

Sales commission  

Product X 10% On  1,68,00,000 = (16,80,000)

   Product Y   20% On 28,12,00,000 = (5,62,40,000)

Depreciation  

Machine M = 1,485,000 / 9 = (1,65,000)

Machine N = 656,000/4 = (1,64,000)

Total Operating expenses = ( 7,34,26,900)

Interest expenses = (0)

Earnings before tax = 20,89,73,100

Tax @ 40% = (8,35,89,240)

Net income after tax = 12,53,83,860

3. Retrun on Equity

Formula = Net inocme divided by Share holders equity

Share holders equity = paid up capital + retained earnings

Paid In Capital = 13,000,000

Additional Paid in Capital = 128,000,000

Retained Earnings = 149,000,000

Share holders equity =   29,00,00,000

Retrun on equity = Net income / share holders equity

= 12,53,83,860 / 29,00,00,000

= 43.2%